OpenStudy (anonymous):

What is the Square root of point zero zero four nine? A.0.0007 B.0.007 C.0.07 D.0.7

OpenStudy (anonymous):

\[\sqrt{0.0049}=0.07 \]

OpenStudy (anonymous):

thank you

OpenStudy (anonymous):

You are welcome.

OpenStudy (anonymous):

maybe more obvious if you write \[\sqrt{.0049}=\frac{\sqrt{49}}{\sqrt{10000}}\]